ANALYSIS OF THE VIBRONIC BANDS IN THE EMISSION SPECTRUM OF p-FLUOROBENZYL RADI-CAL

\underline{SANG K. LEE}, Department of Chemistry, Pusan National University, Pusan 609-735, South Korea; TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, Columbus, Ohio 43210.

We have employed the technique of corona excited supersonic expansion to generate the jet cooled p-fluorobenzyl radical from p-fluorotoluene. The vibronic emission spectra of the p-fluorobenzyl radical has been recorded with a Fourier transform spectrometer. A full vibronic analysis of the spectra has been performed, allowing for unambiguous assignments of the ground state vibrational frequencies of the p-fluorobenzyl radical. In addition, we observed several weaker satellite sequence bands near each of the strong vibronic bands. The characteristics of the low-frequency sequence bands have been tested by varying the experimental conditions. The results of the vibronic analysis and a discussion of the satellite bands will be presented.